The polynomial x 3 + 5x 2 - ­57x -­189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in. If the width of the box is 15 in., what are the other two dimensions? ( Hint: The height is greater than the depth.)

Answer by ankor@dixie-net.com(22740)

The polynomial x^3 + 5x^2 - ­57x - 189 expresses the volume, in cubic inches, of a shipping box, and the width is (x+3) in.
If the width of the box is 15 in., what are the other two dimensions?
(Hint: The height is greater than the depth.)
:
x^3 + 5x^2 - ­57x - 189
:
It says the width is (x+3) then it says the width is 15", therefore:
(x+3) = 15
x = 15 - 3
x = 12
:
Use long division, divide polynomial by the width (x+3)
. . . . . . . . . . . x^2 + 2x - 63
. . . . . -----------------------
(x+3)|x^3 + 5x^2 - 57x - 198
:
FOIL x^2 + 2x - 63
(x+9)(x-7)
:
Find dimensions when x = 12
Length: 12 + 9 = 21 in
depth: 12 - 7 = 5 in
width: 12 + 3 = 15 in
:
21 by 15 by 5 inches are the dimensions